OpenStudy (anonymous):

Let f(x)=-4e^(xcosx) f'(x)=? I need help with this problem, not sure if i am supposed to use chain rule or product rule, if so what are some ways I can recognize which rules to use on such problems?

jimthompson5910 (jim_thompson5910):

the chain rule will come first, then you'll use the product rule to derive x*cos(x)

jimthompson5910 (jim_thompson5910):

it might help to replace x*cos(x) with something like w, so w = x*cos(x) f(x) = -4e^(x*cos(x)) f(x) = -4e^(w) f ' (x) = -4e^(w) * dw/dx when you get to dw/dx, you'll use the product rule

jimthompson5910 (jim_thompson5910):

The general rule is to start on the outside and work your way inside (in terms of functions)

OpenStudy (anonymous):

thank you @jim_thompson5910 So when using the chain rule, I take the derivative of outside and multiply it with the derivative of the inside?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

and you can have it nested as deep as you want (you don't necessarily need to have 2 levels deep)

jimthompson5910 (jim_thompson5910):

example: e^(sin(2x)) outermost function: f(x) = e^x derivative: f ' (x) = e^x middle function: g(x) = sin(x) derivative: g ' (x) = cos(x) innermost function: h(x) = 2x derivative: h ' (x) = 2 ------------------------- so... p(x) = f(g(h(x))) = e^(sin(2x)) p ' (x) = f' (g(h(x))) * g ' (h(x)) * h ' (x) p ' (x) = e^(sin(2x)) * cos(2x) * 2 p ' (x) = 2*e^(sin(2x)) * cos(2x) this is a chain rule example 3 levels deep (so to speak). Again, you can go as deep as you want/need to.

OpenStudy (anonymous):

So would the derivative of the outside be -sinx ?

OpenStudy (anonymous):

Got it, thanks for the explanation !! @jim_thompson5910